# Navigating the Future of Federated Recommendation Systems with Foundation Models

### v1

@@ VI Future Directions
This section discusses applications that highlight the great potential of generative models in federated recommender systems. These applications not only have the potential to improve recommendation quality, but also to provide personalised services while protecting user privacy. As technology continues to advance, we can expect generative models to play an increasingly important role in future FRSs.
VI-A Personalized Content Generation
Generative models such as VAEs and GANs can be used to generate personalized user profiles within federated recommendation systems. These models can capture latent features of user preferences and generate recommendations that reflect these characteristics, providing more accurate personalized suggestions.
VI-B Privacy Protection
Privacy is paramount in federated recommendation systems. Generative models can protect user data privacy by generating representations of user preferences locally and then only sending these representations to a central server for recommendations, without disclosing raw user data.
VI-C Data Augmentation
Diffusion models and large language models can be used to generate additional training data, which is particularly useful in federated learning environments where data sparsity is often an issue. By generating synthetic data, the diversity of the training set can be increased, enhancing the performance of the recommendation system.
VI-D Cross-Domain Recommendations
Federated recommendation systems allow multiple organizations to share in the training of a recommendation model without sharing user data. Generative models can be used within this framework to generate cross-domain recommendations, sharing generated latent user features across different organizations rather than raw data.
VI-E Cold Start Problem Solution
For new users or new items, traditional recommendation systems face the so-called ”cold start” problem. Generative models can predict the latent features of these new entities and generate initial recommendations, quickly initiating the recommendation process.
VI-F Enhanced Recommendation Explainability
Large language models can be used to generate explanations for recommendations, helping users understand why a particular piece of content is recommended. This is crucial for increasing user trust and acceptance of recommendation systems.
VI-G Real-Time Recommendations
Generative models can quickly produce recommendations, which is useful for recommendation scenarios that require real-time feedback, such as news recommendations or real-time event recommendations.
VI-H Reduction of Recommendation Bias
By generating a diverse range of recommendations, generative models can help reduce biases in recommendation systems, providing more equitable recommendation outcomes.
VI-I System Robustness Enhancement
In federated recommendation systems, generative models can be used to simulate various user behaviors and preferences, thereby testing and improving the robustness of the recommendation system.
VI-J Multimodal Recommendations
Combining multiple types of data such as images, text, and audio, generative models can be used to create multimodal recommendations, such as recommending travel destinations that combine images with textual descriptions.
VII Applications
VII-A eHealthcare Systems
eHealthcare systems leverage modern technology to enhance the quality and efficiency of medical services, such as remote diagnosis and patient monitoring. Federated recommendation systems allow different medical institutions to share insights from patient data without sharing the data itself, which is crucial for protecting sensitive health information. Generative models use this data to provide doctors and patients with personalized treatment plans and health advice, improving diagnostic accuracy and treatment outcomes. Thus, federated recommendation systems combined with generative models can integrate patient health data across institutions to create complex treatment pathways and simulate disease progression, enabling doctors to better predict patients’ responses to specific treatments and provide more precise medical services.
VII-B Industrial Applications
In manufacturing, automation and intelligent technologies significantly enhance production efficiency and safety. Federated recommendation systems foster technological collaboration between different enterprises by sharing non-sensitive data, accelerating product design and optimization processes. This not only shortens product development cycles but also protects corporate trade secrets. Generative models can design new mechanical parts or product prototypes by simulating different design variables, quickly proposing multiple design solutions; they can also help predict machine failures, perform maintenance in advance, and reduce downtime, thereby enhancing production efficiency. Federated recommendation systems combined with generative models enable the sharing of improvements and innovations while maintaining data privacy between companies.
VII-C Consumer Services
Analyzing consumer behavior helps businesses deeply understand customer needs. Federated recommendation systems can integrate data from different service providers while protecting user privacy, achieving personalized services. Generative models provide personalized recommendations based on this data, enhancing user experience. For example, in travel recommendation systems, models generate customized travel itineraries based on users’ historical preferences. By simulating different consumer behavior patterns, businesses can predict market trends, adjust service strategies, and enhance competitiveness. Federated recommendation systems combined with generative models analyze user behavior data from different service providers to generate service plans that meet individual needs, promoting personalized service development.
VII-D Retail and E-commerce
Applying recommendation systems to online shopping and retail provides a personalized shopping experience. Federated recommendation systems in the retail and e-commerce sectors analyze consumer shopping habits to generate personalized product recommendations, increasing users’ purchasing desire and satisfaction. Retailers and e-commerce platforms can use generative models to predict consumer buying behavior, thereby offering more accurate product recommendations, optimizing inventory, and enhancing sales efficiency. Federated recommendation systems combined with generative models allow retailers and e-commerce platforms to consider market conditions comprehensively, predicting consumer buying behavior and providing more precise product recommendations, optimizing inventory, and enhancing sales efficiency.
VII-E Finance and Banking
In the financial sector, recommendation systems use data analysis and prediction models to optimize financial services, applicable for credit assessments, stock market analysis, etc., offering clients tailored financial products. Federated recommendation systems ensure transaction data privacy while offering clients tailored financial products. Generative models can simulate market behaviors, predicting stock prices or financial product returns. These models help financial institutions formulate more accurate investment strategies while reducing risks. In finance, federated recommendation systems combined with generative models help financial institutions protect client privacy while simulating market changes and user behavior, offering advice on investment products and credit strategies, thereby enhancing clients’ investment returns and satisfaction.
VII-F Smart Cities
In smart city applications, systems need to process and analyze large-scale urban data. From optimizing traffic flow to devising public safety strategies, federated recommendation systems can help city planners use resident data without directly sharing it to recommend strategies for improving public services and infrastructure, enhancing urban management efficiency and improving residents’ quality of life. Additionally, using generative models to simulate the effects of different strategies can find optimal solutions for managing and allocating urban resources. For example, by simulating traffic flow, models can predict congestion and propose solutions, improving urban traffic efficiency. Federated recommendation systems combined with generative models can help city planners optimize resource allocation and management efficiency without sharing resident data.
VII-G Social Networking
Social platforms facilitate interpersonal communication and information sharing online, using recommendation systems to suggest content to users, such as friends or topics of interest, to increase user engagement. Federated recommendation systems can enhance user engagement and satisfaction while protecting user privacy. Generative models can generate personalized content recommendations, enhancing user engagement and satisfaction while ensuring the protection of user privacy, increasing platform attractiveness and user retention.
VII-H Education
In the field of education, using recommendation systems can improve teaching methods by making appropriate recommendations based on students’ progress and abilities, providing a personalized learning experience. Generative models can generate personalized learning materials and courses based on students’ progress and interests. These customized learning resources can enhance students’ learning efficiency and interest. Federated recommendation systems combined with generative models can generate personalized learning paths based on students’ feedback and learning history. By simulating teaching scenarios and student learning processes, these systems can provide resources that better meet individual learning needs.
VIII Resources
VIII-A Datasets
@@
In FRS, the datasets used can be categorized into two types based on user feedback: explicit feedback and implicit feedback, as shown in Fig. 7. Explicit feedback includes direct responses from users about their preferences. This typically includes ratings (such as a 1-5 scale), user comments, and like/dislike statements. Such data is clear and provides straightforward insights into user preferences, making it highly valuable for training recommendation models. Common datasets with explicit feedback include the following shown in Table ?????????I:
@@ VIII-B Metrics
For predicting how well a system can estimate user preferences, there are several metrices measuring the prediction errors to judge the accuracy of the predictions, such as Mean Absolute Error (MAE), Mean Squared Error (MSE) and Root Mean Squared Error (RMSE). When it comes to classifying items, i.e., determining whether a user will like a product or not, we look at metrics such as Precision, Recall, Hit Ratio (HR), F1 Score, Accuracy and AUC. These tell us how correctly the RS is classifying items, and the F1 score helps us balance the Precision and Recall.
Then to measure the item ranking ability, which is about listing recommendations in the right order, Average Precision (AP) and Mean Average Precision (MAP) are key to this, as they assess the quality of the order of recommendations. Metrics such as Mean Reciprocal Rank (MRR), Normalized Mutual Rank (NMR) and Normalised Discounted Cumulative Gain (NDCG) also contribute by assessing how well the top recommended items are ranked.
On a broader scale, we consider recommendation-centric metrics such as Diversity, which ensures that a variety of items are suggested, and Coverage, which measures how many items from the catalogue are recommended. There are also some user-centric metrics include Novelty, which measures how new or surprising the recommendations are, and Degree of Agreement (DOA), which quantifies the level of concordance between the ranking of items produced by a recommendation system and the ranking preferred by the user. Moreover, business metrics such as Click-Through Rate (CTR) are critical to assessing the system’s impact on user engagement and the company’s bottom line, and Conversion Rate (CVR) measures how efficient an algorithm is at providing recommendations that lead to user purchases. There also are some metrics for measuring other functionality. For example, Gini Index evaluates the fairness of recommendation distribution, with lower values indicating more equitable distribution across items.
Furthermore, in the generative recommendation scenario, FMs such as DMs and LLMs can generate items that have never appeared in historical data and recommend them to users. In this case, how to evaluate the generative recommendation capability of these generative FMs remains an open question.
IX Conclusion

### v2

@@ VI Future Directions
The emerging field of applying FM to FRS presents both opportunities and challenges. The combination of the powerful representation capabilities of FM and the data privacy protection advantages of FRS can enhance the performance of recommendation services and address some of the challenges that traditional FRS methods struggle to overcome. This section discusses some key future directions for FRS with FM, aiming to provide valuable guidance for future research and applications. These directions can offer higher quality personalized recommendation services while protecting user privacy. As technology continues to advance, it is reasonable to posit that FM will play an increasingly important role in future FRS.
VI-A Data Augmentation
FMs, especially generative FMs such as LLMs and DMs, have the ability to generate training data needed for downstream tasks such as text and image generation [150, 151]. They can be used to generate more user interaction data, alleviating data scarcity problems. In FRS, data from different participants may be unbalanced or insufficient. Using FM for data augmentation can effectively increase the variety and amount of training data.
Through data augmentation, FM can help build more comprehensive user profiles and item characteristics, thereby improving the personalization and effectiveness of RS. For example, virtual interaction records of users or additional item descriptions can be generated to better capture user interests and preferences [152]. However, as mentioned earlier, the generated data may have quality issues that could degrade the performance of the RS and introduce biases and errors. Therefore, it is critical to perform rigorous validation and cleaning of the generated data to ensure that the synthetic data is statistically consistent with the real data and to reduce the noise introduced by the synthetic data [132]. Similarly, FM should be combined with privacy-preserving techniques, such as differential privacy, to ensure that no sensitive user information is revealed during the generation of synthetic data. In addition, FM should be used to generate diverse data to cover different scenarios and user behaviors, reducing the risk of bias and over-fitting [153].
VI-B Cold-Start Recommendation
The core of the cold start problem is data sparsity, meaning that new users and new items lack sufficient historical interaction data [154]. This makes it difficult for the model to capture users’ true interests and preferences, which affects the accuracy and effectiveness of recommendations. Because FM has been pre-trained on large amounts of textual data, it has learned rich semantic information that helps generate high-quality representations of users and items. Therefore, FM has strong zero-shot and few-shot learning capabilities, allowing it to provide effective recommendations for cold start users and items even without sufficient interaction data [155, 41]. In addition, FM can transfer knowledge learned in one domain to another through techniques such as fine-tuning and prompt learning. This means that even in cold start situations, the model can use previously learned knowledge to make recommendations, thereby improving the performance of the recommendation system.
It is important to note that in a federated environment, collecting and using enough information to solve the cold start problem while protecting user privacy is a significant challenge. Ensuring the privacy and security of user data is a fundamental requirement for any FRS. In addition, multi-task learning [156] can improve the adaptability and generalization of basic models across different tasks and domains. For example, training on both recommendation and classification tasks simultaneously enables the model to make effective recommendations in more scenarios.
VI-C Multi-Modal Recommendation
FM has the ability to process and understand different modalities of data, such as text, images, audio, and video [18]. In FRS, the integration of multi-modal data enables FM to build more comprehensive and multi-dimensional user profiles, thereby improving the personalization of recommendations [157, 158]. These rich user profiles can more accurately reflect the user’s interests and preferences, helping to deliver recommendations that better meet the user’s needs.
It is important to note that a major challenge of multi-modal data is its heterogeneity, meaning that different modalities of data have different characteristics and structures [159]. For example, text data is sequential, while image data consists of two-dimensional pixels. This heterogeneity increases the complexity of data fusion and processing. Therefore, the development of unified data representation methods that map different modalities of data into a common representation space is essential for model processing and integration.
In a federated environment, special attention must be paid to the privacy of multi-modal data. Multi-modal data may contain more sensitive information, such as the user’s voice or video, which requires more stringent privacy measures to effectively handle and protect this data. In addition, the processing and integration of multi-modal data requires significant computational resources and time. The training and inference processes of FM on multi-modal data can be more complex and time-consuming, making it critical to improve computational efficiency.
VI-D Real-Time Recommendations
Real-time recommendation is a critical task in FRS, which aim to dynamically deliver personalized content based on real-time user behavior and contextual data [160]. This recommendation model emphasizes immediate processing and response to data, ensuring that users receive the most relevant recommendations in the shortest possible time. FM excels at understanding contextual information, which enhances their performance in processing user queries, item descriptions, and other textual data to improve recommendation accuracy and relevance.
However, real-time recommendations require processing and responding to user requests in extremely short time frames, which places high demands on computational resources and system latency. FM typically requires significant computational resources, which can lead to increased latency and negatively impact the user experience. In addition, limitations on the length of the input context can cause performance issues when dealing with long sequences of user behavior or a large number of candidate items.
In real-time recommendations, FM can also reflect biases from the training data in its recommendations, leading to unfair biases. For example, popular items may be preferentially recommended due to their high frequency in the training data, overshadowing long-tail items [136]. Therefore, it is crucial to explore efficient model compression and acceleration techniques, such as distillation [161] and pruning [162], to reduce computational complexity and latency. Introducing strategies such as sliding windows [163] can manage and optimize the input of contextual information, thereby improving the model’s performance in handling long behavioral sequences. Incremental learning methods [164, 165]can also be used to update the FM, and model distillation techniques can reduce model size, thereby lowering communication overhead.
VI-E Enhanced Recommendation Explainability
The federated nature and recommendation characteristics of FRS require transparent explanations of recommendation results to users. This transparency helps users understand the logic and reasoning behind the recommendations, thereby increasing their trust and satisfaction. Language FMs, pre-trained on large amounts of text data, have strong natural language generation capabilities. They can generate clear and coherent textual explanations while referencing relevant background knowledge, helping users to understand the reasons behind recommendations [166, 167]. For example, models such as ChatGPT can generate detailed recommendation explanations that explain why a particular item is suitable for the user [168]. However, generating detailed recommendation explanations often requires significant computational resources [169]. This is especially true for LLMs, which can increase latency and negatively impact the user experience. In addition, these LLMs may exhibit biases present in the pre-training data, resulting in unfair tendencies in the generated explanations, such as gender or race biases. Therefore, it is important to apply model compression and optimization techniques to reduce the computational burden of FM, and to design debiasing methods and fairness constraints to minimize biases in recommendation explanations. Furthermore, it is crucial to incorporate user feedback mechanisms to continuously improve the quality of recommendation explanations. By leveraging user feedback, strategies for generating recommendation explanations can be adapted and optimized, thereby increasing user satisfaction and trust.
VI-F Adavanced Metrics
Traditional evaluation metrics, which will be detailed in VII-B, are designed for specific tasks in recommendation scenarios, such as rating prediction or item ranking. However, the strength of FM in RS lies in its powerful generative capabilities, which allow it to generate items that have never appeared in historical data and recommend them to users. In addition, FM can generate explanations and user interaction content, which requires more complex and comprehensive evaluation criteria. In this context, how to evaluate the recommendation capabilities of these systems remains an open question. Therefore, FRS with FM requires new evaluation metrics tailored to the specific applications of FM in RS, especially in generative recommendation.
VII Resources
VII-A Datasets
@@
In FRS, the datasets used can be categorized into two types based on user feedback: explicit feedback and implicit feedback, as shown in Fig. 7. Explicit feedback includes direct responses from users about their preferences. This typically includes ratings (such as a 1-5 scale), user comments, and like/dislike statements. Such data is clear and provides straightforward insights into user preferences, making it highly valuable for training recommendation models. Common datasets with explicit feedback include the following shown in Table I:
@@ VII-B Metrics
For predicting how well a system can estimate user preferences, there are several metrices measuring the prediction errors to judge the accuracy of the predictions [185], such as Mean Absolute Error (MAE), Mean Squared Error (MSE) and Root Mean Squared Error (RMSE). When it comes to classifying items, i.e., determining whether a user will like a product or not, we look at metrics such as Precision, Recall, Hit Ratio (HR), F1 Score, Accuracy and AUC. These tell us how correctly the RS is classifying items, and the F1 score helps us balance the Precision and Recall.
Then to measure the item ranking ability, which is about listing recommendations in the right order, Average Precision (AP) and Mean Average Precision (MAP) are key to this, as they assess the quality of the order of recommendations. Metrics such as Mean Reciprocal Rank (MRR) [186], Normalized Mutual Rank (NMR) [187] and Normalised Discounted Cumulative Gain (NDCG) [188] also contribute by assessing how well the top recommended items are ranked.
On a broader scale, we consider recommendation-centric metrics such as Diversity, which ensures that a variety of items are suggested, and Coverage, which measures how many items from the catalogue are recommended. There are also some user-centric metrics include Novelty, which measures how new or surprising the recommendations are, and Degree of Agreement (DOA) [189], which quantifies the level of concordance between the ranking of items produced by a recommendation system and the ranking preferred by the user. Moreover, business metrics such as Click-Through Rate (CTR) [190] are critical to assessing the system’s impact on user engagement and the company’s bottom line, and Conversion Rate (CVR) [190] measures how efficient an algorithm is at providing recommendations that lead to user purchases. There also are some metrics for measuring other functionality. For example, Gini Index [191] evaluates the fairness of recommendation distribution, with lower values indicating more equitable distribution across items.
VIII Conclusion